50 grams of ice-cold water (0 °C) is heated to boiling (100 °C). The specific heat capacity of water is 4.18 J /g°C. How much energy did the water absorb?


1
Expert's answer
2021-11-13T01:33:25-0500

m= 50 g

"Q = mcΔT \\

Q = 50 \times 4.18 \times (100 -0) \\

Q = 20900 \;J = 20.9 \;kJ"